OpenStudy (anonymous):

Solve x + 3 = 11. A. –8 B. 8 C. –14 D. 14

OpenStudy (anonymous):

b

OpenStudy (the_fizicx99):

subtract 3 from both sides. x = 8

OpenStudy (anonymous):

Solve 13 – x = 4. A. –17 B. 9 C. –9 D. 17

OpenStudy (anonymous):

transfer 3 from lhs to rhs then x=11-3=8

OpenStudy (the_fizicx99):

@abhi250592 you can get suspended by giving out the answers.

OpenStudy (anonymous):

Solve 13 – x = 4. A. –17 B. 9 C. –9 D. 17

OpenStudy (the_fizicx99):

B, do the same subtract. (inverse)

OpenStudy (anonymous):

Solve 5x = –25. A. –15 B. –4 C. –25 D. –5

OpenStudy (beccaboo333):

one question per post @pauljohn

OpenStudy (anonymous):

Solve 5x = –25. A. –15 B. –4 C. –25 D. –5

OpenStudy (beccaboo333):

it's against OpenStudy policy to post more than one question in a single post.

OpenStudy (anonymous):

Solve 5x = –25. A. –15 B. –4 C. –25 D. –5

OpenStudy (beccaboo333):

You need to stop posting your test questions. One per post.